The radioactive decay rate of a radioactive element is found to be 103 disintegration / second at a certain time. If the half life of the element is one second, the decay rate after one second is . . . . . . . . and after three second is . . . . . . . . . 

Grade:upto college level

1 Answers

Deepak Patra
9 years ago
A = A0 (1/2)n where A0 = Initial activity = 1000 dps (given )
A = Activity after n half lives
At t = 1, n = 1 ∴ A = 1000 (1/2)1 = 500 dps
At t = 3, n = 3 ∴ A = 1000 ( ½ )3 = 125 dps